A total distance of travel of 1000 km breaks every 100 km of a pattern of East- South- East- South- continuously up to the end. The time of travel was recorded 5 hours. (a) Find the average velocity of the travel. (b) if the helicopter flies from the same origin, what will be it's average velocity in an hour travel?
1
Expert's answer
2021-07-15T02:48:59-0400

Part a

Here The distance is broken every 100 km in the south direction then next to east it is repeated every 100 km 

So the We calculate the displacement for the first 100 km south and 100 km of the east.  

The displacement "d_1 = \sqrt{100^2+100^2}=141.42 km"

So the total displacement is D = 5 x 141.42 = 707.1 km 

Part b

f helicopter flies from the same origin at a 1000 Km distance in 1 hour 

The average velocity of the helicopter 

"V = \frac{707.1}{1}=707.1 Km/hour \\
V = \frac{707.1×5}{18}\\
V = 196.42m/s"